# C2P: Featuring Large Language Models with Causal Reasoning

### v2

@@ 3 Developing Causal Chain of Prompting
To develop C2P framework, we aim to extract the adjacency matrix of variables in a given premise. The adjacency matrix acts as equivalent to and instead of the causal DAG in Pearl’s work leading to answering the causal reasoning question. In developing the Causal Chain of Prompting, we use a few steps similar to the PC algorithm for the given premise.
3.1 Causal Chain of Prompting (C2P)
@@
The C2P frameworks consist of 5 main sub-tasks as follows, for learning and reasoning cause and effect relation for the given premise;
  1. 1.
    Prompting to extract the random variables in the provided data.
  2. 2.
@@
    Prompting to extract all the cause and effect relations along with all conditional and unconditional relations among the random variables specifically mentioned in the given premise.
  3. 3.
@@
    Prompting to create an initial adjacency matrix where all elements are 1, except for the diagonal elements and the elements corresponding to the cause-and-effect relationships specifically mentioned in the given premise (extracted in sub-task 2).
  4. 4.
  5. 5.
    Prompting for cause-and-effect questions or hypotheses.
Each sub-task in C2P includes multiple steps and overall, to execute the frameworks, 9 main prompts have to be completed. The exact prompts for all the steps are provided in the Appendix, A.1.
3.2 Few-shot learning agenda
@@
LLMs have powerful zero-shot capabilities, yet they might struggle with complex tasks with insufficient examples in the training process. In these situations, few-shot learning is a versatile and efficient technique for in-context learning that can be used to quickly adapt LLMs to new tasks and significantly enhance their performance (Min et al., 2022; Touvron et al., 2023). This approach involves providing several examples with desired answers, which condition the LLMs to produce the correct response for new instances with similar patterns. The few-shot learning process of C2P is similar to the previous subsection, with the prompts being based on the more abstract prompts presented there (Due to the token limitation). The prompts and an example of the given story are included in the Appendix, A.2. Depending on the token limitation of employed LLMs, the number of given examples (shots) varies. Comparing the token limitation of GPTs, we use GPT-4 Turbo, which has a 30000 token limit.
@@
Steps from other constraint-based methods can be used instead or along with some steps of the PC algorithm in C2P methods, comprehensively discussed in (Glymour et al., 2019). For instance, steps in the FCI method can be used in cases where the causal sufficiency assumption is violated, i.e., latent variables and selection bias may be present (Spirtes et al., 1995). Additionally, it is important that the classifiers proposed by Ceraolo et al. (2024) that formalize the definition of causal questions and establish a taxonomy for finer-grained classification can be used before our framework to identify the causal question as a prerequisite of our method.
4 Results of the experiment for featuring LLMs with Causal Reasoning
@@ 4.1 Experimental setup
The experimental setup is divided into two data groups: synthetic dataset and realistic scenarios.
In a first synthetic dataset, and similar to the experimental setup of (Jin et al., 2023b), we set up the following list of LLMs for the experiments on the CORR2CAUSE dataset. Various types of LLMs are tested in (Jin et al., 2023b) and we only include powerful LLMs in our experiment to avoid reproducing the same results. To test existing LLMs, we first include two BERT-based NLI models in the transformers library (Wolf et al., 2020): BART (Lewis et al., 2019), DistilBART (Shleifer & Rush, 2020). We also evaluate the recent, more efficient models, LLaMa3-8B (Touvron et al., 2023) and LLaMa3-70B (Taori et al., 2023). We assess the GPT-3.5 (i.e., ChatGPT), and the latest GPT-4 (OpenAI, 2023) by April 2023, using the OpenAI API (https://openai.com/api/) with temperature 0. We employed the most recent GPTO as well. Then, we assess C2P on a proposed synthetic dataset, similar to the one in (Jin et al., 2023a), but more complex and with less information provided on the premises. Then we test the reliability of C2P in different numbers of nodes. Due to the limitation of tokens for different versions of GPTs, we employed GP4-Turbo which has a 30000 maximum token limit. We selected 100 samples at random from the corr2cause dataset for our experiment, ensuring a balanced distribution with 50 "Yes" and 50 "No" answers. In the corr2cause dataset, the majority of responses are "No", as a result, they used the F1 score as the primary metric for accuracy. Our experiment design enhances both the realism and comprehensibility of the results. These samples explore reasoning within three causal scenarios: direct cause-and-effect relationships (Fig. 2 i), indirect cause-and-effect relationships (Fig. 2 ii), and the presence of an effect due to two causes (Fig. 2 iii).
@@
To evaluate the proposed framework on natural stories, we utilized the state-of-the-art large language model, GPT-4, which excels at story generation. We crafted detailed instructions in the prompts and generated approximately 20 stories for our case study in fields such as healthcare and medicine, economics, education, social sciences, environmental science, and marketing, all emphasizing the importance of causality. This process is similar to the one presented in (Jin et al., 2023b), but our generated stories are more complex. Our approach can be tested using the examples of the Simpson’s Paradox discussed in (Pearl & Mackenzie, 2018). However, since these examples are already included in the training data of current LLMs, the models simply repeat the correct answers based on that training data, similar to the parrot study in (Zečević et al., 2023). Consequently, our generated natural stories replicate these examples in a manner that the LLM cannot address within its existing training data.
Additionally, we assess the coevolution of supermassive black holes (SMBHs) and their host galaxies using our proposed framework, replicating the results provided in (Pasquato et al., 2023).
4.2 Evaluation of the C2P on synthetic dataset
Results of the C2P on CORR2CAUSE dataset: In Table LABEL:table4, we show the performance of the more powerful LLMs in the cause-effect task reported in (Jin et al., 2023b) compared to the results of the C2P framework.
Table 1: Main results: Comparison of using C2P frameworks compared to the existing LLMs
@@ 4.2 Evaluation of the C2P on synthetic dataset
LLaMa3-8B 48.97 51.06 47.05 51
LLaMa3-70B 70.8 60.7 56.6 60
@@
GPT-3.5 48.8 55 44 54
GPT-4 48.4 53.3 48 53
GPT-4O 57.7 59.5 56 59
C2P on GPT-4 Turbo 92.78 95.74 90 93
Few-shot learned GPT-4 Turbo 81.06 82.97 79.24 81
According to Table LABEL:table4, performing pure causal reasoning tasks presents a significant challenge for all existing LLMs. Notably, several models perform worse than a random guess, indicating a complete failure in this specific task of pure causal reason. Nevertheless, applying our proposed framework to GPT-4 Turbo has boosted at least 33% in the model’s ability to respond to causal inquiries, suggesting a substantial improvement over the existing capabilities of LLMs. It is important to note that the primary factor contributing to the discrepancy between the LLaMa and GPT results and those reported in (Jin et al., 2023b) stems from the distribution of the provided premise. This table indicates that when reasoning questions are posed to LLMs, the responses slightly tend towards the "No" answer.
Robustness analysis: To assess the robustness of the proposed prompts, Table 2 shows the computed accuracy computed in each sub-task of implementing C2P on GPT-4 for different numbers of variables in the given premise.
Table 2: Accuracy by Number of Variables and sub-tasks
@@
First sub-task Second sub-task Third sub-task Fourth sub-task Fifth sub-task
3 variables 100% 100% 100% 100% 99%
4 variables 100% 100% 100% 98% 83.2%
5 variables 100% 100% 100% 88.3% 75.2%
6 variables 100% 100% 100% 79.1% 70.1%
4.3 Evaluation of the C2P on natural stories
To evaluate the proposed framework on natural stories, we utilize the same process presented in (Jin et al., 2023b) and generate premises. The result of applying the C2P framework, both step-by-step and few-shot learned, is provided in Table 3 and compared to the result of GPTs. This table clearly demonstrates that the existing method responds randomly to reasoning prompts. It is important to note that even when GPTs provide the correct final answer, the reasoning process leading to that answer is still incorrect. In contrast, GPT-4 Turbo, enhanced with step-by-step prompting of the C2P framework and few-shot learned C2P, boosts the accuracy of these models in similar scenarios involving natural stories.
Table 3: Result of applying the C2P framework and GPTs, TP: True Positives, FP: False Positives, TN: true negatives, FN: False Negatives

### v4

@@ 3 The Causal Chain of Prompting
To develop our framework, our aim is to extract the adjacency matrix of variables on a given premise. The adjacency matrix acts as equivalent to and instead of the causal DAG in Pearl’s work, leading to answering the causal reasoning question. In developing the Causal Chain of Prompting, we use a few steps similar to the PC algorithm for the given premise.
3.1 Causal Chain of Prompting ($\text{C}^{2}\text{P}$)
@@
The $\text{C}^{2}\text{P}$ framework consists of five main subtasks as follows, for learning and reasoning about cause and effect relations for the given premise:
  • Subtask 1: Prompting to extract the random variables in the provided data.
@@
  • Subtask 2: Prompting to extract all the cause-and-effect relations along with all conditional and unconditional relations among the random variables specifically mentioned in the given premise.
@@
  • Subtask 3: Prompting to create an initial adjacency matrix where all elements are 1, except for the diagonal elements and those corresponding to the cause-and-effect relationships specifically mentioned in the given premise (extracted in subtask 2).
  • Subtask 4: Prompting of conditional and unconditional independence evaluation and identification of colliders to extract the causal PDAG.
  • Subtask 5: Prompting for cause-and-effect questions or hypotheses.
Each subtask in $\text{C}^{2}\text{P}$ can include one or multiple steps (prompts). In general, to execute the framework, 9 main steps must be completed. Subtask 1 is completed in Step 1. Subtask 2 is accomplished with Step 2. Subtask 3 is achieved with Step 3. To perform Subtask 4, 5 steps must be applied. Step 4 first eliminates all unconditional independencies achieved in Subtask 2 and Step 5 then removes all conditional independencies extracted in Subtask 2. Step 6 identifies potential nodes that can act as colliders. Step 7 confirms whether the nodes identified in Step 6 are colliders. Step 8 updates the adjacency matrix, resulting in the final adjacency matrix of a causal structure. To perform Subtask 5, the causal question is asked in Step 9. The exact prompts for all the steps are provided in Appendix A.3 and code repository.
3.2 Few-shot learning
@@
LLMs have powerful zero-shot capabilities, yet they struggle with complex tasks because of their engineering designs and insufficient examples in the training process. In such situations, few-shot learning is a versatile and efficient technique for in-context learning, which can be used to quickly adapt LLMs to new tasks and significantly enhance their performance (Min et al., 2022; Touvron et al., 2023). This approach involves providing several examples with desired answers to condition the LLMs to produce correct responses for new instances with similar patterns. The few-shot learning process of $\text{C}^{2}\text{P}$, as described in the previous subsection, is based on more abstract prompts (due to token limitations). The prompts and an example of the given story are included in Appendix A.4. Depending on the token limitations of the employed LLMs, the number of examples (shots) may vary. For GPT-4 Turbo, which has a 30,000 token limit, we were able to include ten examples. For LLaMA 3.1, we used the same examples in the few-shot learning process as well.
@@
It is important to note that steps from other constraint-based methods can either replace or complement the steps of the PC algorithm in $\text{C}^{2}\text{P}$ methods, as comprehensively discussed in (Glymour et al., 2019). For instance, the steps in the FCI method can be used in cases where the causal sufficiency assumption is violated, i.e., where latent variables and selection bias may be present (Spirtes et al., 1995). Additionally, the classifiers proposed by Ceraolo et al. (2024) formalize the definition of causal questions and establish a taxonomy for finer-grained classification. These classifiers can be used before our framework to identify the causal question as a prerequisite of our method.
4 Experiments
4.1 Datasets
The experiments are divided into two data groups: a synthetic dataset and a real-world scenarios dataset.
The CORR2CAUSE dataset:
Introduced in (Jin et al., 2023b), the CORR2CAUSE dataset serves as a benchmark to assess the ability of LLMs to respond to reasoning queries. Tab. LABEL:table01 shows the details of the samples in this dataset. The process of creating CORR2CAUSE is as follows: First, select the number $N$ of variables (Step 1) and generate all unique DAGs with $N$ nodes (Step 2). Next, gather all d-separation sets from these graphs to identify MECs (Step 3). In Step 4, formalize the data by associating each MEC with its corresponding causal graphs. For each MEC, construct a correlation statement based on the statistical relations within the MEC, hypothesize a causal relationship between two variables, and assign a validity "Yes" if the hypothesis holds for all causal graphs in the MEC, or "No" if the hypothesis does not necessarily apply to all graphs in the MEC. Finally, introduce the verbalization process. An example of a premise and its corresponding hypothesis in the CORR2CAUSE dataset is as follows:
Premise: Suppose that there is a closed system of 3 variables, A, B and C. All statistical relations among these 3 variables are as follows: A correlates with C. B correlates with C. However, A is independent of B. Hypothesis: A directly affects C.
Table 1: Information on CORR2CAUSE samples, different variable numbers, and responses.
Scenario Varible 3 varibles 4 varibles 5 varibles 6 varibles Sum
CORR2CAUSE test data
All samples 45 72 518 522 1157
Samples with "No" answers 30 57 449 402 938
Samples with "Yes" answers 15 15 69 120 338
CORR2CAUSE train data
All samples 0 576 7524 197634 205734
Samples with "No" answers 0 529 6542 160615 167686
Samples with "Yes" answers 0 47 982 37019 38048
The Natural Stories dataset:
The Natural Stories dataset is also introduced in (Jin et al., 2023b) to assess the reasoning capabilities of LLMs in realistic scenarios. This dataset builds upon the CORR2CAUSE dataset as a foundation for future extensions in various contexts, such as instantiating variables with real-world phenomena and placing the narratives in more natural settings. For instance, the rule "correlation does not imply causation" can be illustrated using ice cream sales and swimming pool attendance as variables, by arguing that ice cream sales do not necessarily impact swimming pool attendance, as both could be influenced by a third factor, such as hot weather. The Natural Stories data presented in (Jin et al., 2023b) is not open-source. However, generating such data is straightforward. By providing an example from the CORR2CAUSE dataset along with the corresponding Natural Stories created by a human or existing example on the web, GPT-4 can generate realistic, everyday narratives. For more detailed examples of generating such stories, refer to the code directory accompanying this paper. A natural story generated by GPT-4, inspired by the previous example, could be:
Premise: Let’s consider three factors: eating junk food, obesity, and watching television. There is a correlation between eating junk food and obesity, and between watching television and obesity. However, eating junk food and watching television are independent from each other. Hypothesis: Eating junk food directly affects obesity.
More details and examples are provided in the code repository of this study.
An example on supermassive black holes.
The data and results on the coevolution of supermassive black holes (SMBHs) and their host galaxies, presented in (Pasquato et al., 2023), are used as a real-world example in its verbalized form. The verbalized information is as follows:
Premise: In the (co)evolution of supermassive black holes (SMBHs) and their host galaxies, the data supports the following observations. Existing studies show that with a change in central density, there is no significant change in bulge stellar mass. However, there is a decrease in black hole mass with higher central density. Additionally, when black hole mass is fixed, central density and velocity dispersion change simultaneously. Conversely, higher velocity dispersion or effective radius results in lower bulge stellar mass, while velocity dispersion and effective radius do not change simultaneously. Hypothesis: Does central density affect black hole mass?
@@ 4.2 Experimental setup
To test existing LLMs on the first synthetic data, we first include two BERT-based NLI models in the transformers library (Wolf et al., 2020): BART (Lewis et al., 2019), DistilBART (Shleifer & Rush, 2020). We evaluate LLaMA3-8B (Touvron et al., 2023) and LLaMA3-70B (Taori et al., 2023). We also evaluate the latest, more efficient models, LLaMA 3.1 8B and LLaMA 3.1 70B released in late September 2024. We assess the GPT-3.5 (i.e., ChatGPT), and the GPT-4 (Achiam et al., 2023), using the OpenAI API (https://openai.com/api/). We also used the latest GPT-4o model, released on October 2, 2024. Then, we evaluate the reliability of $\text{C}^{2}\text{P}$ on various queries, each with a different number of random variables. Due to the limitation of tokens for different versions of GPTs, we employed GP4-Turbo which has a 30000 maximum token limit. We also used $\text{C}^{2}\text{P}$ with LLaMA 3.1 70B as well. In all of the models, we set the temperature to 0. Note that various types of LLMs were tested in (Jin et al., 2023b), and the results showed that existing LLMs perform worse than random, completely random, or only slightly better than random in responding to such queries. From the models tested in (Jin et al., 2023b), we selected those with the highest accuracies to avoid duplicating the same results and to stay focused on the main objective of this study.
According to Tab. LABEL:table01, the majority of responses in CORR2CAUSE dataset are "No" ($81.5\%$). Additionally, most samples involved in reasoning have six variables ($96.06\%$), a challenging task even for humans. As a result of these biases, the F1 score is used as the primary metric for evaluating accuracy in CORR2CAUSE simulations. We resolve these biases, by using 420 samples. First, to ensure a balanced distribution of samples, we select a population with 60 "Yes" and 60 "No" answers and an equal number of samples for each number of variables. This sample size is determined based on two main considerations. First, to ensure significant improvements in both step-by-step prompting and few-shot experiments, we applied the sample size formula for comparing two proportions as described by (Chow et al., 2017) and the minimum required sample size is 117. Additionally, to address the limited number of "Yes" responses (only 15 samples), we selected 30 samples for each variable number to minimize bias, resulting in a total of 120 samples. To ensure that the improvement of the results remains significant for the higher number of samples, we use 300 samples with 5 random variables with equal "Yes" and "No" answers. This experimental design enhances the realism and comprehensibility of the computational metrics, providing a more accurate reflection of the model’s performance. As a result, all four key metrics—F1 score, precision, recall, and accuracy—are practical, each offering a unique perspective on the model’s capabilities. These samples explore reasoning within three causal scenarios: direct cause-and-effect relationships (Fig. 2 i), indirect cause-and-effect relationships (Fig. 2 ii), and the presence of an effect due to two causes (Fig. 2 iii).
@@
To evaluate the proposed framework for natural stories, we used GPT-4, which excels at story generation. We crafted instructions in the prompts and generated 30 stories for our case study in fields such as healthcare and medicine, economics, social sciences, environmental science, and marketing, all highlighting the importance of causality. This process is similar to the one presented in (Jin et al., 2023b). Our approach can be tested using the examples of Simpson’s paradox discussed in (Pearl & Mackenzie, 2018). However, since these examples are already included in the training data of current LLMs, the models simply repeat the correct answers based on that training data, similar to the parrot study in (Zečević et al., 2023). Consequently, our generated natural stories replicate these examples in a manner that the LLM cannot address within its existing training data. We aimed to demonstrate how the symbolic expressions in the CORR2CAUSE dataset can affect the reasoning of our proposed framework.
Finally, we assess the coevolution of supermassive black holes (SMBHs) and their host galaxies using our proposed framework, replicating the results from (Pasquato et al., 2023). The goal in (Pasquato et al., 2023) was to extract a PDAG of SMBHs based on numerical data and then infer causal reasoning questions from the graph. We use verbalized information on SMBHs to evaluate whether the LLM, enhanced with $\text{C}^{2}\text{P}$, can answer reasoning questions such as, "Does central density affect black hole mass?"
It is important to note that Chain-of-Thought (CoT) prompting is implemented in most existing LLMs, enhancing their reasoning accuracy, as discussed in (Chung et al., 2024). Consequently, in our experiments, LLMs are prompted to think step by step when responding to reasoning questions, ensuring that the CoT mechanism is activated.
4.3 Evaluation of the $\text{C}^{2}\text{P}$ on synthetic dataset
Results of the $\text{C}^{2}\text{P}$ on CORR2CAUSE dataset: In Tab. 2, we show the performance of LLMs in the cause-effect task with and without employment of $\text{C}^{2}\text{P}$ framework.
Table 2: Comparison of applying $\text{C}^{2}\text{P}$ frameworks in LLMs compared to the existing LLMs with CoT
@@ 4.3 Evaluation of the $\text{C}^{2}\text{P}$ on synthetic dataset
LLaMA 3-8B 43.37 48.6 43.15 47.5
LLaMA 3-70B 49.41 53.84 62.23 55.77
LLaMA 3.1-8B 53.96 45.94 61.41 46.66
LLaMA 3.1-70B 57.14 48.64 49.15 48.07
$\text{C}^{2}\text{P}$ with LLaMA 3.1-70B 81.63 83.3 83.3 81.66
$\text{C}^{2}\text{P}$ Few-shot learned LLaMA 3.1-70B 76 79.1 73.07 76.66
@@
GPT-3.5 47.5 56.2 43.2 52.5
GPT-4 Turbo 51.9 51.92 45 54.2
GPT-4 50.2 54.1 47.1 55
GPT-4O 59.86 61.46 58.33 60.83
$\text{C}^{2}\text{P}$ with GPT-4 Turbo 91.72 93.47 89.2 91.66
$\text{C}^{2}\text{P}$ Few-shot learned GPT-4 Turbo 81.6 79.21 77.21 79.28
According to Tab. 2, performing causal reasoning tasks remains a significant challenge for existing LLMs, even when prompted to think step by step, similar to (Wei et al., 2022b). These results show that even more updated models can sometimes perform worse than older ones in some metrics, for instance, LlaMa 3.1 acts worse than LlaMa 3. Based on the responses presented in the code repository for this study, model hallucination is one of the major factors contributing to this poor performance, aligning with the findings in Jin et al. (2023b). Tab. 2 shows that by applying 5 consecutive tasks of $\text{C}^{2}\text{P}$ on GPT-4 Turbo and LlaMa 3.1, the reasoning accuracy improved by over $30.7\%$ and $25.9\%$, respectively, compared to the corresponding LLMs without $\text{C}^{2}\text{P}$ on a synthetic benchmark dataset. Then, using few-shot learning of GPT-4 Turbo and LLaMA 3.1 with $\text{C}^{2}\text{P}$, reasoning accuracy increased by over $20.05\%$ and $20.89\%$, respectively, with as few as ten examples, compared to the corresponding LLMs without $\text{C}^{2}\text{P}$ on the same dataset. To ensure that the improvements are significant in both step-by-step prompting and few-shot experiments, we applied the sample size formula for comparing two proportions from (Chow et al., 2017). This formula indicates that even with a sample size of 117, the observed differences are statistically significant we assume a confidence level of $99\%$ and $80\%$ power. It is important to note that the primary factor contributing to the discrepancy between the results of LLaMA and GPT, and those reported in (Jin et al., 2023b), is the distribution of the provided premise. This table shows that when reasoning questions are posed to GPT models, their responses tend to slightly favor the "No" answer. Interestingly, LLaMA models slightly tend to answer "Yes" to reasoning questions more frequently. It is important to note that even when LLMs provide the correct final answer, the reasoning process that leads to that answer can still be incorrect, as illustrated in A.2, Tab. LABEL:tab:_wrong_and_right, for the results of GPT-4. This highlights the randomness of their responses more clearly which is also investigated and shown in Bao et al. (2024) as well. Other prompts that can execute the algorithm can be used in place of ours, as long as they perform the same subtasks. This supports the idea presented in the Lu et al. (2024). Furthermore, this becomes particularly evident in the few-shot learning process, where providing more examples has a significant impact; even more abstract prompts can achieve the same or even better results.
Robustness and reliability analysis: To assess the robustness of the proposed prompts for each of the 5 subtasks, Tab. 3 shows the computed accuracy computed in each subtask of implementing $\text{C}^{2}\text{P}$ (subsection 3.1) on GPT-4 Turbo for different numbers of variables in the given premise.
Table 3: Accuracy by number of variables and subtasks
@@
First subtask Second subtask Third subtask Fourth subtask Fifth subtask
3 variables 100% 100% 100% 99.12% 98.7%
4 variables 100% 100% 100% 97.5% 84.1%
5 variables 100% 100% 100% 87.5% 75%
6 variables 100% 100% 100% 78.3% 70%
The sample size formula for comparing two proportions indicates that fewer than 120 samples are sufficient to demonstrate a significant improvement in reasoning performance when using $\text{C}^{2}\text{P}$. Figure 3 shows results for datasets with sample sizes ranging from 30 to 300. This figure demonstrates that LLMs, such as GPT-4 Turbo, consistently achieve higher performance across all four metrics when $\text{C}^{2}\text{P}$ is applied, compared to scenarios where $\text{C}^{2}\text{P}$ is not used.
Figure 3: The accuracy metrics for different numbers of samples with 3 different methods.
According to this figure, the $\text{C}^{2}\text{P}$ framework enhances all metrics across varying sample sizes.
4.4 Evaluation of the $\text{C}^{2}\text{P}$ on natural stories
The results of applying the $\text{C}^{2}\text{P}$ framework, both step-by-step and few-shot learned for GPT-4 Turbo and LLaMA 3.1, are provided in Tab. 4 and compared to the results of GPTs and LLaMA 3.1. This table, along with the highlighted sections, clearly demonstrates that the existing methods respond randomly to reasoning prompts. In contrast, GPT-4 Turbo and LLaMA 3.1 with step-by-step prompting of the $\text{C}^{2}\text{P}$ framework and few-shot learned $\text{C}^{2}\text{P}$, significantly improve accuracy in scenarios involving natural stories.
Table 4: Comparison of applying $\text{C}^{2}\text{P}$ frameworks with existing LLMs with CoT in responding to queries within natural stories, TP: True Positives, FP: False Positives, TN: true negatives, FN: False Negatives.